# High-frequency tails in spectral densities

### v1

@@ Abstract
Recent developments in numerically exact quantum dynamics methods have brought the dream of calculating the dynamics of chemically complex open systems closer to reality. Path-integral-based methods, hierarchical equations of motion (HEOM) and quantum analog simulators all require the spectral density of the environment to describe its effect on the system. Here we find that the rate of slow population relaxation is sensitive to the precise functional form used to describe the spectral density peaks. This finding highlights yet another challenge to obtaining accurate spectral densities. We give a simple recipe to adjust the results of quantum simulation for this difference assuming both the simulator and the target spectral densities are known.
I Introduction
@@
Accurate calculations of quantum dynamics of open quantum systems with chemically complex environments would advance our understanding of many problems of interest in chemistry, biology, and quantum information science. Tuning of the coherence times of molecular qubits,[1] learning from efficient energy transfer in photosynthesis,[2] and in silico design of molecular engines[3] - these are just select examples of the many new possibilities that would open up. Due to the inherent complexity of the open quantum dynamics, there isn’t a single universal approach to solve it. Instead, a great variety of methods has been developed, each of them with a different regime of applicability.
@@
The main challenge in open quantum dynamics is to accurately describe the effect of a large environment on a small system of interest. The unitary dynamics of system plus environment are intractable owing to the large (possibly macroscopic) environment. Various approximate methods have been developed, that utilize convenient but often severe approximations: Markovian, secular, finite order perturbation theory, semiclassical, etc. These methods are much less demanding in terms of computational cost, but the validity of various approximations needs to be verified. We do not consider our findings in the context of these approximate methods. Historically, two main strategies to make the problem tractable without introduction of uncontrolled approximations emerged; we will refer to these two strategies as "unitary" and "reduced" for brevity. In the unitary approach, finite number of degrees of freedom of the environment are included and the total dynamics (of both system and truncated environment) is unitary. Multiconfigurational time-dependent Hartree (MCTDH),[4, 5] and its multilayer extension,[6] the density matrix renormalization group (DMRG),[7] the time-evolving density matrix using the orthogonal polynomials algorithm (TEDOPA),[8] the time-dependent Davydov ansatz,[9] and the effective-mode (EM) approach[10] all utilize this strategy. Only finite number of environment modes can be included, meaning that the overall dynamics is reversible. Thus, the system cannot reach thermal state even in principle and instead will experience recurrences in long-time dynamics. On the other hand, these methods are applicable to "tough" cases, such as anharmonic environment modes,[11] failure of the Born-Oppenheimer approximation,[12] or strong,[13] possibly nonlinear[5] coupling to the environment.
@@
Hierarchical equations of motion (HEOM)[14] and its variants,[15, 16, 17] as well as real-time path integral (PI) methods[18, 19, 20, 21, 22, 23, 24] utilize the reduced strategy to make open quantum dynamics numerically tractable. "Reduced" here refers to the fact that only system dynamics is followed explicitly; the effect of the environment is captured implicitly by introducing a bath of large (possibly uncountably infinite) number of harmonic degrees of freedom, each bilinearly coupled to the system. The effect of this harmonic bath on the system dynamics on one hand can be captured exactly within an effective description,[25, 26] and on the other hand it can mimic[27, 28] the effect of a complex environment through Gaussian response. For example, recently, the relaxation in bacterial light harvesting chromophore has been simulated[29] using the small matrix decomposition of the path integral algorithm.[21] Similarly, HEOM was used to model the FMO complex.[30] While extensions to nonlinear system-bath couplings[31, 32] as well as approaches that utilize (anharmonic) spin baths[33, 34] have been developed, we will only consider the simplest case of linear coupling to harmonic bath, as this is the setting used in the vast majority of numerically exact calculations to date.
@@
Recent advances on the intersection between quantum information science and theoretical chemistry opened up a radically new set of approaches to obtain accurate open quantum dynamics – using digital quantum computation[35, 36] or analog quantum simulation.[37, 38] The former has many of the strengths and weaknesses of the unitary methods, since quantum computers are naturally suited to describe unitary quantum dynamics. In contrast, quantum analog simulators can be used to realize open system dynamics with macroscopically large environments,[38] which groups them together with the reduced methods.
@@
Both PI and HEOM methods are numerically exact, meaning the dynamics of the system can be described as accurately as needed by tightening convergence parameters appropriately. [^1] Of course, in practice even within the domain of applicability of PI and HEOM methods we are limited by available computational resources, as well as by numerical instabilities. Assuming the dynamics is converged, the model we formulate to describe a process is the only remaining source of discrepancies between a numerically exact simulation and physical reality. In this paper we analyze the importance of faithful representation of a structured bath, illustrating our findings using the HEOM simulations.[17] However, we emphasize that our findings extend to any reduced approach (all HEOM variants, all real time PI methods, and quantum analog simulators with continuous baths).
@@
The properties of any harmonic bath bilinearly coupled to the system are fully captured[25] by the bath spectral density (SD) defined in the frequency domain as
@@
with each bath mode $\omega_{j}$ characterized by the system-bath coupling constant $g_{j}$ and $\hbar$ denoting the reduced Planck’s constant. Note that there are alternative definitions of the spectral density in the literature with different prefactors[26] or different powers of frequency, for example Ref. [40] defines spectral density as $j(\omega)=J(\omega)/\omega^{2}$, which is more convenient for the interpretation of the optical properties.[41]
The spectral density enters the equations of motion of the open quantum system via the bath correlation function (BCF) also known as bath response function,[25]
@@
where as usual $\beta=(k_{B}T)$ is the inverse temperature multiplied by the Boltzmann constant, and $f_{BE}$ denotes the Bose-Einstein distribution $f_{BE}(x)=(1-e^{-x})^{-1}$.
Obtaining an accurate SD for a system interacting with a structured environment is highly nontrivial. Molecular dynamics (MD) or the hybrid quantum mechanics/molecular mechanics (QM/MM) methods can be used to calculate classical (i.e., real only) BCF and the SD can be constructed from it. [42, 43, 44] Experimental data can also inform construction of the SDfor a pure dephasing process. For instance, fluorescence line narrowing spectra[45, 46] or resonance Raman scattering spectra.[47] However, both experimental and theoretical approaches yield limited accuracy, such that the precise functional form of the SD that determines its decay rate at high frequency cannot be determined at present. Moreover, most implementations of HEOM variants require that the BCF is written as a finite sum of exponentials,[17, 15, 48, 16, 49]
@@
with complex prefactors $c_{j}$ and complex $\Omega_{j}=\omega_{j}+i\gamma_{j}$, whose real and imaginary parts are related to the central frequency and broadening of the peak in spectral density. We note that the decompositions using other types of functions are possible within HEOM as well, for example as a sum of Bessel functions of first kind.[50] We will only consider exponential BCFs here.
Therefore, the choice of the functional form of SD in open dynamics simulations is frequently motivated by either the experimental constraints (for instance, when using RLC (resistor-inductor-capacitor) circuits to mimic bath degrees of freedom in quantum analog simulators)[38] or by the numerical efficiency considerations (as in HEOM variants where the Drude-Lorenz form is particularly efficient),[51, 41] or by the exact results in limiting cases (for example underdamped Brownian peaks which describe one vibrational mode coupled to the system as well as to a secondary harmonic bath).[52, 51, 41] Additionally, other numerically exact methods like quantum-classical path integral[22, 23] require the discretization of the bath, which necessarily implies a high-frequency cutoff and can obscure the functional form of the peak
In what follows, we will analyze the effect of using different functional forms to describe the features (i.e., peaks) of a given SD. The condition of Eq. (3) requires that each peak of the spectral density is represented with a fit function that has a small (but nonzero) number of first order (i.e., simple) poles in the lower half of the complex plane (see section II.2). We will also limit ourselves to only considering odd functions of frequency, such that the resulting spectral density is odd by construction, as required by Eq. (1). While these conditions might seem restrictive at first sight, we note that two of the most commonly used functional forms to model the spectral density (Drude-Lorenz and underdamped Brownian oscillators) satisfy them.[41, 51]
There are several excellent and comprehensive analyses of the impact of low-frequency end of the spectrum (i.e., $\lim_{\omega\rightarrow 0^{+}}J(\omega)\propto\omega^{s}$) on the dynamics of spin-boson and displaced oscillator models,[26] where $s=1$ (ohmic), $s<1$ (subohmic), and $s>1$ (superohmic) cases exhibit qualitatively distinct dynamics. By contrast, we have not found studies that analyze in detail the high-frequency end of the spectrum, which is the focus of this work.
The paper is organized as follows. We begin section II with a brief reminder of open quantum dynamics, then in section II.1 we describe the HEOM approach that allowed us to include chemically complex spectral density at tractable cost; we discuss representations of the spectral density peaks in section II.2. Finally, we detail the model system parameters in section II.3. We analyze the impact of the SD tails on the dephasing (section III.1) and population relaxation (section III.2) of a realistic system (electronic excitation of thymine-like molecule in water) and wrap up in section III.3 with the way to connect simulations with different SD basis functions valid in the limit where dephasing and population relaxation are decoupled by separation of timescales.
II Methods
@@
II.1 HEOM
We describe the dynamics of the system with the HEOM.[17] Initially the system is assumed to be in a separable state with the bath at inverse temperature $\beta$, such that the total density matrix at time 0 is
@@ II.1 HEOM
where $Z_{b}=\operatorname{Tr}_{b}{e^{-\beta\hat{H}_{b}}}$ is the bath partition function and $\operatorname{Tr}_{b}$ denotes a partial trace, taken over all bath degrees of freedom. While it is possible to also treat initially entangled states by e.g. shifting the definition of initial time, we only consider initial product states as shown in Eq. (5).
To simplify notation, we omit the hats over density operators and reserve $\rho$ with appropriate subscripts to denote a (possibly reduced) density matrix throughout.
The total (unitary) dynamics of system and environment is unitary and is generated by Eq. (4). However, if only the system dynamics is of interest, the environment degrees of freedom can be traced over to yield the reduced density matrix
@@
This reduced density matrix has nonunitary dynamics given by[15]
@@
and determines the spectral density (Eq. (1)) that appropriately describes the environment’s influence on the system dynamics.
II.2 Spectral density decompositions
We consider a situation where the spectral density of the environment is known from experiment, simulation or the combination of the two and moreover this known spectral density consists of a broad low-frequency ($<300$ cm-1) feature and a finite number of sharp peaks in the $300-4000$ cm-1 frequency range. This is a typical situation for molecules in solution, where the electronic energy levels are affected by a number vibrational peaks as well as the low-frequency collective motion of the solvent.
We therefore approximate the full spectral density as a sum of discrete peaks, each characterized by the peak position (frequency) $\omega_{k}$, peak width (broadening) $\gamma_{k}$ and peak intensity (reorganization energy) $\lambda_{k}$ resulting in the functional form $J_{k}(\omega;\omega_{k},\gamma_{k},\lambda_{k})$. Rather than concentrating on the low-frequency behavior of $J_{k}$ (i.e., ohmic, subohmic, superohmic), we will focus our attention on the decay of high-frequency tails at the opposite end of the spectrum. Fig. 1(a) shows a single peak at $1500$ cm-1 for each of the functional forms considered in this study. The five functional forms with identical peak parameters yield near-perfect agreement in the vicinity of the peak (Fig.  1b,c). The last two panels highlight the low- (Fig. 1d) and high- (Fig. 1e) frequency behavior of the different functional forms. Note that panel (d) is a log-log plot, while (e) has linear x-axis and logarithmic y-axis.
The displaced Drude oscillator has Lorenzian shape
@@ II.2 Spectral density decompositions
The sum in the first line of Eq. (11) ensures the odd symmetry of the fit function. The integral of $J_{k}^{Dr}/\omega$ over the full range of frequency is the reorganization energy of the $k$th peak, $\lambda_{k}$. Note that the ubiquitous Drude-Lorenz (DL) form $J^{DL}(\omega)=\frac{2\lambda\gamma\omega}{\pi(\omega^{2}+\gamma^{2})}$ is a special case of the displaced Drude oscillator (Eq. (11)) when the displacement $\omega_{k}$ is equal to 0. It corresponds to one summand in BCF (Eq. (3)), that does not oscillate ($\omega_{k}=0$) but has a pure exponential decay with rate $\gamma_{k}$, a special case for which HEOM simulations are particularly efficient.
Carrying out similar analysis for the underdamped Brownian oscillator (UBO) yields the functional form given in Eq. (14). Note that the UBO SD has Ohmic behavior at low frequency (Eq. (15)) with a coefficient that’s double that of the displaced Drude oscillator (compare with Eq. (12)). However, high-frequency tails fall off faster than the Drude case, i.e., as $\omega^{-3}$ vs $\omega^{-1}$.
@@
The UBO SD is physically motivated, as it corresponds to a situation where the system is directly and linearly coupled to a single (nuclear) mode of frequency $\omega_{k}$, which in turn is coupled to a bath of modes undergoing Brownian motion with friction $\gamma_{k}$.[53]
The recent proposal for the quantum analog simulator device[38] utilizes RLC circuits connected to the gate defined quantum dots to simulate a two-level system linearly coupled to a bosonic bath. Harmonic oscillators are mechanical analogs of the LC (inductor-capacitor) circuits and the presence of the resistive element "R" introduces broadening of discrete peaks to yield a spectral density of a functional form very similar to the UBO (compare Eq. (17) to Eq. (14)), but the low frequency behavior is superohmic (Eq. (18)) and the high-frequency tail falls off as slowly as the Drude peak, i.e., as $\omega^{-1}$ but with double the prefactor (compare Eq. (19) to Eq. (13)).
@@
Each of the three SD functional forms (Eqs. (11), (14), (17)) affects the dynamics of an open quantum system by adding two summands to the BCF (Eq. (3)), that decay exponentially with rate $\gamma_{k}$ and oscillate with frequency $\omega_{k}$. Taking $\gamma_{k}$ and $\omega_{k}$ to be the same in all three cases thus presents a natural point of comparison between the three functional forms. In such a comparison the difference between the three SD functional forms is fully contained within BCF coefficients $c_{j}$ (Eq. (3)).
In addition to the spectral density basis functions described above, we also tested more complicated functions, that can be seen as generalizations of the UBO modes.[41] Each of these 2-peak functions corresponds to two pairs of terms in the BCF, that oscillate at frequencies $\{\omega_{k1},\omega_{k2}\}$ respectively and exponentially decay with rates $\{\gamma_{k1},\gamma_{k2}\}$ respectively:
$J_{k}^{2,n}(\omega)$ $=$ $\frac{\Lambda_{n,k}\omega^{n}}{\prod\limits_{i=\{1,2\}}\left((\omega-\omega_{ki})^{2}+\gamma_{ki}^{2}\right)\left((\omega+\omega_{ki})^{2}+\gamma_{ki}^{2}\right)}$ (20)
$J_{k,0}^{2,n}(\omega)$ $=$ $\frac{\Lambda_{n,k}\omega^{n}}{\prod\limits_{i=\{1,2\}}(\omega_{ki}^{2}+\gamma_{ki}^{2})^{2}}+\mathcal{O}(\omega^{n+2})$ (21)
@@
where the frequency scaling at the low- and high-frequency ends are $\omega^{n}$ and $\omega^{-(8-n)}$ respectively for $n\in\{1,3,5,7\}$. Here we employ the first two of the four possible basis functions of this form and refer to them as "Ohmic with 2 peaks" ($n=1$) and "Superohmic with 2 peaks" ($n=3$). The corresponding normalization constants $\Lambda_{n,k}$ are given in the Appendix A. These functions can have second order poles, but for the present discussion we will choose parameters that ensure the poles are simple.
@@
II.3 Model system
For simplicity we consider a two level system model of a molecule in aqueous solution. We use electronic parameters as well as the SD obtained for thymine in aqueous solution.[47] However, we emphasize that this is not done to faithfully describe the dynamics of thymine, but rather to make the model parameters realistic. If accurate description of thymine was our goal instead, we would need to account for the conical intersection, that leads to fast ($\sim 400$ fs) population relaxation in the lab.[54]
The parameters for model Hamiltonian (Eq. (4)) are obtained as follows: The energy gap between two levels $\Omega=35,650$ cm-1 ($4.42$ eV) is extracted from experimenal absorption spectrum of thymine in aqueous solution.[55] We construct the bath using the recently reported SD parameters obtained for thymine in aqueous solution based on resonance Raman experiments.[47] While these parameters are obtained in the pure dephasing limit, we wish to explore the regime where bath-assisted population relaxation is possible as well. We therefore couple the bath via both $\sigma_{z}$ with $\alpha_{z}=0.37$ and $\sigma_{x}$ with $\alpha_{x}=0.34$. The precise choice of coefficients is somewhat arbitrary, but to allow for both dephasing and population relaxation they are chosen to be of comparable magnitude. We use the frequencies and reorganization energies obtained in Ref. 47 (see Table 1).
The resonance Raman experiments do not inform the widths of the peaks $\gamma_{k}$, so we have to make a reasonable choice. The simplest option would be to pick a single width (say 10 $\mathrm{cm}^{-1}$) for all the peaks in the spectral density. This works well for the displaced Drude oscillator, underdamped Brownian, and the RLC circuit functional forms, where intensity of each peak is controlled by the corresponding reorganization energy parameter $\lambda_{k}$. However, for ohmic and superohmic 2-peak functions reorganization energy parameter $\Lambda_{2p,k}$ sets the total peak intensity of a pair of peaks (Eqs. 25-26). The peak widths in Table 1 are chosen (as described in Appendix A) to ensure that the Ohmic 2-peak functions yield individual peak intensities shown in Table 1.
Table 1: Spectral density parameters for our model system. The first two columns are values for thymine obtained from Ref. 47, the third is constructed based on the behavior of Ohmic 2-peak function as described in the main text.

### v2

@@ Abstract
Recent advances in numerically exact quantum dynamics methods have brought the dream of accurately modeling the dynamics of chemically complex open systems within reach. Path-integral-based methods, hierarchical equations of motion (HEOM) and quantum analog simulators all require the spectral density (SD) of the environment to describe its effect on the system. Here we focus on the decoherence dynamics of electronically excited species in solution in the common case where nonradiative electronic relaxation dominates and is much slower than electronic dephasing. We show that the computed relaxation rate is highly sensitive to the choice of SD representation –– such as the Drude-Lorentz or Brownian modes – or strategy used to capture the main SD features, even when early-times dephasing dynamics remains robust. The key reason is that electronic relaxation is dominated by the resonant contribution from the high-frequency tails of the SD, which are orders of magnitude weaker than the main features of the SD and can vary significantly between strategies. This finding highlights an important, yet overlooked, numerical challenge: obtaining an accurate spectral density requires capturing its structure over several orders of magnitude to ensure correct decoherence dynamics at both early and late times. To address this, we provide a simple transformation that recovers the correct relaxation rates in quantum simulations constrained by algorithmic or physical limitations on the shape of the SD. Our findings enable comparison of different numerically exact simulation methods and expand the capabilities of analog simulations of open quantum dynamics.
[…]
SI
1 Introduction
@@
Accurate calculations of the quantum dynamics of open quantum systems with chemically complex environments would advance our understanding of many problems of interest in chemistry, biology, and quantum information science. Tuning of the coherence times of molecular qubits,1 learning from efficient energy transfer in photosynthesis,2 and in silico design of molecular engines3 - these are just select examples of the many new possibilities that would open up. Due to the inherent complexity of the open quantum dynamics, there isn’t a single universal approach to solve it. Instead, a great variety of methods has been developed, each of them with a different regime of applicability.
@@
The main challenge in open quantum dynamics is to accurately describe the effect of a large environment on a small system of interest. The unitary dynamics of system plus environment is intractable owing to the large (possibly macroscopic) environment. Historically, two main strategies to make the problem tractable without the introduction of uncontrolled approximations have emerged;we will refer to these two strategies as "unitary" and "reduced" for brevity. In the unitary approach, finite number of degrees of freedom of the environment are included and the total dynamics (of both system and truncated environment) is unitary. Multiconfigurational time-dependent Hartree (MCTDH),4, 5 and its multilayer extension,6 the density matrix renormalization group (DMRG),7 the time-evolving density matrix using the orthogonal polynomials algorithm (TEDOPA),8 the time-dependent Davydov ansatz,9 and the effective-mode (EM) approach10 all utilize this strategy. Only a finite number of environment modes can be included, meaning that the overall dynamics is reversible. Thus, the system cannot reach the thermal state even in principle and instead will experience recurrences in long-time dynamics. On the other hand, these methods are applicable to "tough" cases, such as anharmonic environment modes,11 failure of the Born-Oppenheimer approximation,12 or strong,13 possibly nonlinear5 coupling to the environment.
@@
Hierarchical equations of motion (HEOM)14 and its variants15, 16, 17 as well as real-time path integral (PI) methods18, 19, 20, 21, 22, 23, 24 utilize the reduced strategy to make open quantum dynamics numerically tractable. "Reduced" here refers to the fact that only system dynamics is followed explicitly; the effect of the environment is captured implicitly by introducing a bath of large (possibly uncountably infinite) number of harmonic degrees of freedom, each bilinearly coupled to the system. The effect of this harmonic bath on the system dynamics on one hand can be captured exactly within an effective description,25, 26 and on the other hand it can mimic27, 28 the effect of a complex environment through Gaussian response. For example, recently, the relaxation in bacterial light harvesting chromophore has been simulated29 using the small matrix decomposition of the path integral algorithm.21 Similarly, HEOM was used to model the FMO complex.30
@@
Recent advances at the intersection of quantum information science and theoretical chemistry opened up a radically new set of approaches to obtain accurate open quantum dynamics – using digital quantum computation31, 32 or analog quantum simulation, where the dynamics of interest is mapped onto a highly controllable experimental system.33, 34, 35, 36 The former has many of the strengths and weaknesses of the unitary methods, since quantum computers are naturally suited to describe unitary quantum dynamics. In contrast, quantum analog simulators can be used to realize open system dynamics with macroscopically large environments,34, 35 which groups them together with the reduced methods.
@@
Both the PI and HEOM methods are numerically exact, meaning the dynamics of the system can be described as accurately as needed by tightening convergence parameters appropriately. Assuming the dynamics is converged, the model we formulate to describe a process is the only remaining source of discrepancies between a numerically exact simulation and physical reality. In this paper, we analyze the importance of faithful representation of a structured bath, illustrating our findings using HEOM simulations.17 However, we emphasize that our findings extend to any reduced approach (all HEOM variants, all real-time PI methods, and quantum analog simulators of open quantum dynamics).
@@
The properties of any harmonic bath bilinearly coupled to the system are fully captured25 by the bath spectral density (SD) defined in the frequency domain as
@@
with each bath mode $\omega_{j}$ characterized by the system-bath coupling constant $g_{j}$ and $\hbar$ denoting the reduced Planck’s constant. The spectral density enters the equations of motion of the open quantum system via the bath correlation function (BCF) also known as bath response function,25
@@
where as usual $\beta=(k_{B}T)^{-1}$ is the inverse temperature multiplied by the Boltzmann constant, and $n$ denotes the Bose-Einstein distribution $n(x)=(e^{x}-1)^{-1}$.
Obtaining an accurate SD for a system that interacts with a structured environment is highly non-trivial. Molecular dynamics (MD) or hybrid quantum mechanics/molecular mechanics (QM/MM) methods can be used to calculate classical (i.e., real only) BCF and the SD can be constructed from it. 37, 38, 39, 40, 41 Experimental data, such as fluorescence line narrowing spectra42, 40 or resonance Raman scattering spectra,43 can also inform the construction of SD for a pure dephasing process. However, both experimental and theoretical approaches have a limited regime of validity and imperfect accuracy.44 The low frequency end of the spectrum (that is, $\lim_{\omega\rightarrow 0^{+}}J(\omega)\propto\omega^{s}$) is known to sensitively affect the dynamics,26 leading to qualitatively distinct features in the ohmic($s=1$), subohmic ($s<1$) and superohmic ($s>1$) cases. In contrast, the details of the high-frequency end of the spectrum are believed to be largely inconsequential with the notable exception of the polaronic dressing of charge carriers.45
[…]
In this study, we reexamine this assumption with a focus on the dynamics of coherent electronic excitation in condensed phase. The decoherence dynamics typically proceeds in two steps: initial fast loss of phase information, followed by the much slower population relaxation. We aim to accurately capture both the fast and the slow components of the decoherence dynamics. We focus on the HEOM method, requiring that the BCF is written as a finite sum of exponentials,17, 15, 53, 16, 54
@@
with complex prefactors $c_{j}$ and complex $\Omega_{j}=\omega_{j}+i\gamma_{j}$, whose real and imaginary parts are related to the central frequency and broadening of the peak in spectral density. This requirement prevents us from directly analyzing the SD with exponential cutoff. Nonetheless, we emphasize that our conclusions about the impact of the high-frequency tails of the spectral density extends beyond the HEOM method with implications for quantum analog simulation as well as other numerically exact methods that employ alternative high-frequency cutoff strategies, including the exponential cutoff. In what follows, we will analyze the effect of using different functional forms to describe the features (i.e., peaks) of a given SD.
The paper is organized as follows. In Sec. 2.1 we describe the Hamiltonian model and summarize the HEOM method. In Sec. 2.2, we discuss representations of the spectral density peaks and, in Sec. 2.3, we detail the model system parameters. Then, in Sec. 3, we analyze the impact of the SD tails on the dephasing (Sec. 3.1) and population relaxation (Sec. 3.2) of an illustrative model system (electronic excitation of a molecule in water) and suggest the way to connect simulations with different SD basis functions (Sec. 3.3). In Sec. 3.4 we discuss the implications and limitations of our findings. We summarize our observations in Sec. 4.
2 Methods
@@
In Eq. (4) we consider the two-level system for simplicity, but the multi-state generalization is straightforward and the conclusions of our analysis carry over. Note that the assumption of harmonic bath does not restrict our considerations to the chemical environments that are harmonic, since the effects of an arbitrary anharmonic environment can be taken into account within Eq. (4) by using an appropriate spectral density, provided the interaction is described well without going beyond second order in perturbation theory,26 which is expected for the macroscopically large environment in the thermodynamic limit, where interaction is distributed over many degrees of freedom.55
2.1 HEOM
We describe the dynamics of the system with the HEOM.17 Initially the system is assumed to be in a separable state with the bath at inverse temperature $\beta$, such that the total density matrix at time 0 is
@@ 2.1 HEOM
where $Z_{b}=\operatorname{Tr}_{b}{e^{-\beta\hat{H}_{b}}}$ is the bath partition function and $\operatorname{Tr}_{b}$ denotes a partial trace, taken over all bath degrees of freedom. To simplify the notation, we omit the hats over density operators and reserve $\rho$ with appropriate subscripts to denote a (possibly reduced) density matrix throughout.
The total dynamics of the system and environment is unitary and is generated by Eq. (4). However, if only the system dynamics is of interest, the environment degrees of freedom can be traced out to yield the reduced density matrix
@@
This reduced density matrix has nonunitary dynamics given by15
@@
and determines the spectral density [Eq. (1)] that appropriately describes the environment’s influence on the system dynamics.
2.2 Spectral density decompositions
We consider a situation where the spectral density of the environment is known from experiment, simulation, or a combination of the two, and, moreover, this known spectral density consists of a broad low-frequency ($<300$ cm-1) feature and a finite number of sharp peaks in the $300-4000$ cm-1 frequency range. This is a typical situation for molecules in solution, where the electronic energy levels are affected by several vibrational peaks, as well as the low-frequency collective motion of the solvent.
We therefore approximate the full spectral density as a sum of discrete peaks, each characterized by the peak position (frequency) $\omega_{k}$, peak width (broadening) $\gamma_{k}$, and peak intensity (reorganization energy) $\lambda_{k}$, resulting in the functional form $J_{k}(\omega;\omega_{k},\gamma_{k},\lambda_{k})$.
The displaced Drude oscillator has Lorentzian shape
@@ 2.2 Spectral density decompositions
The sum in the first line of Eq. (11) ensures the odd symmetry of the fit function. The integral of $J_{k}^{Dr}/\omega$ over the entire frequency range is the reorganization energy of the $k$th peak, $\lambda_{k}$. Note that the ubiquitous Drude-Lorentz (DL) form $J^{DL}(\omega)=\frac{2\lambda\gamma\omega}{\pi(\omega^{2}+\gamma^{2})}$ is a special case of the displaced Drude oscillator (Eq. (11)) when the displacement $\omega_{k}$ is equal to 0. It corresponds to one summand in BCF [Eq. (3)], which does not oscillate ($\omega_{k}=0$) but has a pure exponential decay with rate $\gamma_{k}$, a special case for which HEOM simulations are particularly efficient.
Performing a similar analysis for the underdamped Brownian oscillator (UBO) yields the functional form given in Eq. (14). Note that the UBO SD has Ohmic behavior at low frequency [Eq. (15)] with a coefficient that is double that of the displaced Drude oscillator [compare with Eq. (12)]. However, high-frequency tails fall off faster than in the Drude case, that is, as $\omega^{-3}$ vs $\omega^{-1}$.
@@
The UBO SD is physically motivated, as it corresponds to a situation where the system is directly and linearly coupled to a single (nuclear) mode of frequency $\omega_{k}$, which in turn is coupled to a bath of modes undergoing Brownian motion with friction $\gamma_{k}$.56
The recent proposal for the quantum analog simulator device34 utilizes RLC circuits connected to the gate defined quantum dots to simulate a two-level system linearly coupled to a bosonic bath. Harmonic oscillators are mechanical analogs of the LC (inductor-capacitor) circuits and the resistive element "R" introduces broadening of discrete peaks to yield a spectral density of a functional form very similar to the UBO [compare Eq. (17) to Eq. (14)], but the low-frequency behavior is superohmic [Eq. (18)] and the high-frequency tail falls off as slowly as the Drude peak, that is, as $\omega^{-1}$ but with double the prefactor [compare Eq. (19) to Eq. (13)].
@@
Each of the three functional forms of the SD [Eqs. (11), (14), (17)] affects the dynamics of an open quantum system by adding two summands to the BCF [Eq. (3)], that decay exponentially with rate $\gamma_{k}$ and oscillate with frequency $\omega_{k}$. Taking $\lambda_{k}$, $\gamma_{k}$, and $\omega_{k}$ to be the same in all three cases presents a natural point of comparison between the three functional forms. In such a comparison the difference between the three SD functional forms is fully contained within BCF coefficients $c_{j}$ [Eq. (3)].
In addition to the spectral density basis functions described above, we also tested more complicated functions that can be seen as generalizations of the UBO modes.49 Each of these 2-peak functions corresponds to two pairs of terms in the BCF, which oscillate at frequencies $\{\omega_{k1},\omega_{k2}\}$ respectively and exponentially decay with rates $\{\gamma_{k1},\gamma_{k2}\}$ respectively:
$J_{k}^{2,n}(\omega)$ $=$ $\Lambda_{n,k}\frac{\omega^{n}}{\prod_{i=\{1,2\}}\left((\omega-\omega_{ki})^{2}+\gamma_{ki}^{2}\right)\left((\omega+\omega_{ki})^{2}+\gamma_{ki}^{2}\right)}$ (20)
$J_{k,0}^{2,n}(\omega)$ $=$ $\Lambda_{n,k}\frac{\omega^{n}}{\prod_{i=\{1,2\}}(\omega_{ki}^{2}+\gamma_{ki}^{2})^{2}}+\mathcal{O}(\omega^{n+2})$ (21)
@@
where the frequency scaling at the low- and high-frequency ends are $\omega^{n}$ and $\omega^{-(8-n)}$ respectively for $n\in\{1,3,5,7\}$. Here we employ the first two of the four possible basis functions of this form and refer to them as "Ohmic with 2 peaks" ($n=1$) and "Superohmic with 2 peaks" ($n=3$). The corresponding normalization constants $\Lambda_{n,k}$ are given in the SI. These functions can have second order poles, but for the present discussion we will choose parameters that ensure the poles are simple.
Another form of the cutoff function – exponential – is often used in path-integral as well as master-equation based open quantum dynamics simulations, but cannot be directly addressed with HEOM since the exponential function has a pole of infinite order, which violates the requirement of Eq. (3). Therefore, we did not include the exponential cutoff function in our numerical simulations. Nonetheless, our analysis is general and has implications for simulations performed with the exponential cutoffs as we discuss in Sec. 3.3
Rather than concentrating on the low-frequency behavior of $J_{k}$ (i.e., ohmic, subohmic, superohmic), we focus our attention on the decay of high-frequency tails at the opposite end of the spectrum. Fig. 1(a) shows a single peak at $1500$ cm-1 for each of the functional forms considered in this study. The five functional forms with identical peak parameters show near perfect agreement in the vicinity of the peak (Fig. 1b,c). The last two panels highlight the low- (Fig. 1d) and high- (Fig. 1e) frequency behavior of the different functional forms. Note that panel (d) is a log-log plot, while panel (e) has a linear x-axis and a logarithmic y-axis.
@@
2.3 Model system
We consider the simplest model to describe molecular decoherence in aqueous solution. The system consists of two electronic levels separated by a frequency $\Omega=35,650$ cm-1 ($4.42$ eV), corresponding to the wavelength of $280~{}nm$57 We construct the bath using the recently reported SD parameters obtained based on resonance Raman spectra of thymine in aqueous solution.43 While these parameters are obtained in the pure dephasing limit, we wish to explore the regime where bath-assisted population relaxation is also possible. We therefore couple the bath via both $\sigma_{z}$ with $\alpha_{z}=0.37$ and $\sigma_{x}$ with $\alpha_{x}=0.34$. The precise choice of coefficients is somewhat arbitrary, but to allow for both dephasing and population relaxation, they are chosen to be of comparable magnitude. We use the frequencies and reorganization energies obtained in Ref. 43 (see Table 1). We emphasize that this paper does not aim to faithfully describe the dynamics of a particular molecule (thymine), but rather to study the general phenomenon with realistic model parameters.
Resonance Raman experiments do not inform the widths of the peaks $\gamma_{k}$, so we have to make a reasonable choice. The simplest option would be to pick a single width (say 10 $\mathrm{cm}^{-1}$) for all peaks in the spectral density. This works well for the displaced Drude oscillator, the underdamped Brownian, and the RLC circuit functional forms, where the intensity of each peak is controlled by the corresponding reorganization energy parameter $\lambda_{k}$. However, for ohmic and superohmic 2-peak functions, the reorganization energy parameter $\Lambda_{2p,k}$ sets the total peak intensity of a pair of peaks [Eqs. (S1)-(S2)]. The peak widths in Table 1 are chosen (as described in the SI) to ensure that the Ohmic 2-peak functions yield the individual peak intensities $\lambda_{k}$ obtained from the resonance Raman. . Note that the peaks in Table 1 are ordered such that neighboring peaks (e.g. 1 and 2) form a single 2-peak function.
Table 1: Spectral density parameters for our model system. The first two columns are obtained from Ref. 43, the third is constructed based on the behavior of Ohmic 2-peak function as described in the SI.